# Residual Connections Harm Abstract Feature Learning in Masked Autoencoders

### v2

@@ 6 Conclusion
Huh et al. [28] raise a key insight in their work – that how a neural network is parameterized matters for fitting the data – and investigate the inductive low-rank bias of stacking more linear layers in a network.
In this work, we observe that the ubiquitous residual network [2] may not be the ideal network parametrization for representation learning and propose a modification of the shortcut path in residual blocks that significantly improves unsupervised representation learning. We explore the connection between our reparameterization of the residual connection and the effective rank of the learned features, finding a correlation between good representations and low-rank representations.
@@ Appendix A Training and Evaluation Details
Our training configurations primarily followed the guidelines established by He et al. [1]. In the ImageNet-1K experiment, our model was trained for 800 epochs, utilizing the AdamW [67] optimizer with a constant weight decay of 5e-2 for a batch size of 1024. We set the maximum learning rate to 6e-4. Initially, the learning rate started at 0 and linearly increased to its maximum over the first 40 epochs, after which it followed a cosine schedule to gradually decrease to zero by the end of the training period. It is worth noting that the learning rate per sample, or effective learning rate, in our setup matched that of He et al. [1], although our maximum learning rate was set lower due to our batch size being a quarter of theirs. We applied random resizing, cropping, and horizontal flipping during training as part of our augmentation scheme. To enhance the quality of the learned representations in most experiments, we employed the normalized pixel loss, as proposed by [1]. For visualization purposes, as shown in Figure 6, we used L2 pixel-wise loss. In the ImageNet-100 experiment, we employed the identical training configuration used in the ImageNet-1K experiments. We train our model with 4 NVIDIA A40 GPUs and a completed trianing usually takes 20 hours on ImageNet-100 and 200 hours on ImageNet-1k.
@@
For the ImageNet-1k dataset, we use the exact same evaluation protocols employed in He et al. [1], which includes random data augmentation.
@@
Figure 5: We present our enhanced UNet Transformer architecture for Masked Auto-encoder. (1) Left: Our customized encoder blocks, equipped with our proposed decay identity shortcuts. (2) Middle: Standard transformer blocks as the decoder blocks. (3) Right: We incorporate the decay identity shortcuts exclusively within the encoder blocks of our UNet transformer and employ standard transformer blocks for the decoder. To support abstract representation learning at the bottleneck, i.e., the last layer of the Encoder 12, we adopt the UNet [25] architecture and create skip connections that transmit every other encoder feature directly to the decoder.
@@ A.3 Modified Architecture
We present a visualization of our UNet transformer design, as outlined in Section  3.2, in Fig. 5. It’s important to note that decayed identity shortcuts are exclusively implemented within the encoder block. Additionally, we establish skip connections from alternating blocks in the encoder to the decoder, following the UNet [68] architecture’s design principles.
@@ B.3 Abstraction and Low-rank in the Supervised Setting
(a) Effective rank of converged supervised ResNet for different depths.
(b) Effective rank of supervised ResNet over training epoch.
Figure 7: We train ResNet models for classification on a small subset of ImageNet and visualize (a) effective rank across different depths at convergence and (b) training dynamics of effective rank over time for various $\alpha_{\rm min}$. In (a) we see that at convergence, our method consistently decreases the feature rank with various depth and, in (b), this pattern is also shown for standard ResNet model at every stage of training.

### v3

@@ 7 Conclusion
Huh et al. (2021) raise a key insight in their work – that how a neural network is parameterized matters for fitting the data – and investigate the inductive low-rank bias of stacking more linear layers in a network. In this work, we observe that the ubiquitous residual network (He et al., 2016) may not be the ideal network parametrization for representation learning and propose a modification of the shortcut path in residual blocks that significantly improves unsupervised representation learning. We explore the connection between our reparameterization of the residual connection and the effective rank of the learned features, finding a correlation between good representations and low-rank representations.
@@ Appendix A Training and Evaluation Details
Our training configurations primarily followed the guidelines established by He et al. (2022). In the ImageNet-1K experiment, our model was trained for 800 epochs, utilizing the AdamW Loshchilov & Hutter (2019) optimizer with a constant weight decay of 5e-2 for a batch size of 1024. We set the maximum learning rate to 6e-4. Initially, the learning rate started at 0 and linearly increased to its maximum over the first 40 epochs, after which it followed a cosine schedule to gradually decrease to zero by the end of the training period. It is worth noting that the learning rate per sample, or effective learning rate, in our setup matched that of He et al. (2022), although our maximum learning rate was set lower due to our batch size being a quarter of theirs. We applied random resizing, cropping, and horizontal flipping during training as part of our augmentation scheme. To enhance the quality of the learned representations in most experiments, we employed the normalized pixel loss, as proposed by He et al. (2022). In the ImageNet-100 experiment, we employed the identical training configuration used in the ImageNet-1K experiments. We train our model with 4 NVIDIA A40 GPUs and a completed trianing usually takes 20 hours on ImageNet-100 and 200 hours on ImageNet-1k.
@@
For the ImageNet-1k dataset, we use the exact same evaluation protocols employed in He et al. (2022), which includes random data augmentation.
@@
Figure 7: We present our enhanced UNet Transformer architecture for Masked Auto-encoder. (1) Left: Our customized encoder blocks, equipped with our proposed decay identity shortcuts. (2) Middle: Standard transformer blocks as the decoder blocks. (3) Right: We incorporate the decay identity shortcuts exclusively within the encoder blocks of our UNet transformer and employ standard transformer blocks for the decoder. To support abstract representation learning at the bottleneck, i.e., the last layer of the Encoder 12, we adopt the UNet Ronneberger et al. (2015) architecture and create skip connections that transmit every other encoder feature directly to the decoder.
@@ A.3 Modified Architecture
We present a visualization of our UNet transformer design, as outlined in Section  3.2, in Fig. 7. It’s important to note that decayed identity shortcuts are exclusively implemented within the encoder block. Additionally, we establish skip connections from alternating blocks in the encoder to the decoder, following the UNet Ronneberger et al. (2015) architecture’s design principles.
@@ B.3 Abstraction and Low-rank in the Supervised Setting
(a) Effective rank of ResNet for different depths at the convergence of the training.
(b) Effective rank of ResNet over training epoch.
Figure 9: Dynamics of the feature rank in the supervised setup. We train ResNet models for a supervised classification task on a small subset of ImageNet. And visualize (a) effective rank across different depths at convergence and (b) training dynamics of effective rank over time for various $\alpha_{\rm min}$. In (a) we see that at convergence, our method consistently decreases the feature rank with various depth and, in (b), this pattern is also shown for standard ResNet model at every stage of training.
@@
Figure 10: Qualitative comparison of images generated by diffusion models. Our method, decayed identity shortcuts with $\alpha_{\rm min}=0.6$, shows improved representation learning and produces higher-quality generated images compared to the baseline, which employs full residual connections ($\alpha_{\rm min}=1.0$).